OpenStudy (anonymous):

Find the circumference of the outer walkway. The radius of the larger circle is 15' + 5' = 20'. r=15

OpenStudy (anonymous):

Believe me if i'm wrong but you take the r multiply it by two to get the diameter which is 30'. Once you do that you're going to want to multiply it by pi, pi = 3.14. You're going to take 30' and multiply that by 3.14. I don't know if i'm correct on this but that's what I learned in my sixth grade class.

OpenStudy (anonymous):

The answer is NOT 3.14, you just have to multiply 30 by 3.14,

OpenStudy (mathmath333):

Circumference \( C=2\pi r\\ where ~~\pi=\dfrac{22}{7}=3.14\)

OpenStudy (mathmath333):

so \(\Large c=2\pi r\\~\\ =2\times3.14\times 15\\~\\ =94.247\)